# Comparative study of the butterfly velocity in holographic QCD models at finite temperature and chemical potential

### v2

@@ 6 Conclusion and Future Outlook
In this work, we conducted a detailed investigation of the butterfly velocity in the deconfined plasma phase of EMD-inspired holographic QCD models. Our analysis focused on four distinct yet physically motivated models that span both top-down and bottom-up approaches within the holographic framework. Specifically, we examined two top-down constructions – the 1RCBH and 2RCBH models – derived from consistent truncations of type IIB supergravity. In addition, we analyzed a semi-analytic bottom-up model based on the potential reconstruction technique, which allows for analytic control over the metric functions and facilitates matching to QCD-like thermodynamics. Finally, we considered a fully numerical bottom-up model, designed to reproduce lattice QCD thermodynamic data at finite temperature and chemical potential, thereby providing a phenomenologically grounded dual description. This diverse set of models enables us to probe the robustness of chaotic transport properties in strongly coupled non-conformal plasmas, particularly in relation to the butterfly velocity and its thermal and chemical potential dependent behavior in QCD.
@@
In each of these holographic models, we computed the butterfly velocity using three distinct methods: entanglement wedge reconstruction, OTOCs, and pole-skipping, and found that all three give the same result. This extends previous results by demonstrating that such an equivalence holds in a variety of holographic QCD models as well. In the first three models, we obtained results analytically, while in the fourth model, we obtained results numerically. We further analyzed the behavior of $v_{B}$ with respect to the chemical potential and temperature and observed a universal trend across models. In particular, $v_{B}$ is found to be decreasing/increasing with the chemical potential/temperature. The same result is also true in the simplest holographic charged plasma, i.e., dual to the Reissner-Nordstrom AdS (RN-AdS) background (see Appendix A). Such universal behavior offers potential signatures of quantum chaos in QCD that could be accessible to experimental observation, especially in light of the recently proposed experiments to measure OTOCs in quantum systems. Such universal chaotic features might also provide interesting insights into QCD.
@@
Another important remark that we would like to make is regarding the behavior of $v_{B}$ with respect to the charge. In this work, we primarily focused on studying the variation of $v_{B}$ with respect to the chemical potential, but one can also study the variation with respect to the charge. On analyzing, one finds that $v_{B}$ monotonically decreases even with the charge for all the models presented in our work. It is interesting to note that such a monotonically decreasing trend with charge was also observed in (hyperscaling-violating) Lifshitz theories [58]. Thus, it would be interesting to investigate if this trend is a universal feature of all holographic theories or all quantum systems in general.
@@
In this work, we restricted our attention to bulk black hole geometries that are isotropic and planar. One can also extend this analysis to black hole geometries that are anisotropic. The work [100] provides an example to compute butterfly velocity in anisotropic backgrounds. Exploring the effect of magnetic fields on the butterfly velocity in holographic QCD models is another interesting avenue. However, since, in most holographic models magnetic field induces an anisotropy, computing the butterfly velocity may not be an easy task. One exception to this is the model discussed in [101]. This model has a $SO(3)$ symmetry in spite of the magnetic field. For this model, one can check, using the methods discussed in this work, that the butterfly velocity monotonically decreases as a function of the magnetic field. It would be interesting to investigate if this decreasing trend is also a feature of other more sophisticated holographic models of magnetized QCD [102, 103].
@@
Before concluding, we would like to make a brief remark on the interplay between the butterfly velocity and phase transitions. In [104], it was conjectured that $v_{B}$ can serve as a probe of phase transitions. That work presented examples within AdS/CMT models supporting this idea. The bottom-up holographic model considered in section 4 exhibits a Hawking-Page phase transition. Notably, we have plotted $v_{B}$ only for the thermodynamically stable branch. If one extends the plot of $v_{B}$ across all branches, a non-monotonic behavior emerges, with the point of non-monotonicity coinciding with the phase transition. This observation is in agreement with the conjecture proposed in [104]. Note that by taking different profiles of the scale factor $A(z)$, one can also construct gravity solutions exhibiting van der Waals type first-order small/large black hole phase transitions with a second-order critical point [52]. It would be interesting to see if information about such transitions and critical points can also be reflected in the butterfly spectrum. There are already interesting suggestions that such small/large phase transitions can be probed by the Lyapunov exponents [105, 106].
@@ Acknowledgements
A.0.1 Background
In this section, we compute $v_{B}$ of the charged plasma corresponding to the Reissner-Nordstrom AdS (RN-AdS) black hole background. The metric of an RN-AdS black hole in five dimensions is
@@ A.0.1 Background
A.0.2 Calculation of $v_{B}$ using entanglement wedge method
Comparison of (A.1) with (2.1) yields the following
@@ A.0.2 Calculation of $v_{B}$ using entanglement wedge method
Substituting the horizon values of the metric functions (and their derivatives) into (2.13), we get the following expression for the butterfly velocity
@@
Figure 9 illustrates the behavior of $v_{B}$ with the RN-AdS background. As shown, the qualitative behavior of $v_{B}$ in this setup closely mirrors the trends observed in the other holographic models discussed in previous sections
A.0.3 Calculation of $v_{B}$ using the OTOC method
@@
A.0.4 Calculation of $v_{B}$ using pole-skipping method
@@
The expression for the butterfly velocity $v_{B}^{2}=\omega_{*}^{2}/q_{*}^{2}$
$v_{B}^{2}=\frac{\omega^{2}}{q^{2}}=\frac{2}{3}\left(1-\frac{Q^{2}}{2}\right)\,.$ (A.11)

### v3

@@ 6 Conclusion and Future Outlook
In this work, we conducted a detailed investigation of the butterfly velocity in the deconfined plasma phase of EMD-inspired holographic QCD models. Our analysis focused on four distinct yet physically motivated models that span both top-down and bottom-up approaches within the holographic framework. Specifically, we examined two top-down constructions – the 1RCBH and 2RCBH models – derived from consistent truncations of type IIB supergravity. In addition, we analyzed a semianalytic bottom-up model based on the potential reconstruction technique, which allows for analytic control over the metric functions and facilitates matching to QCD-like thermodynamics. Finally, we considered a fully numerical bottom-up model, designed to reproduce lattice QCD thermodynamic data at finite temperature and chemical potential, thereby providing a phenomenologically grounded dual description. This diverse set of models enables us to probe the robustness of chaotic transport properties in strongly coupled nonconformal plasmas, particularly in relation to the butterfly velocity and its thermal- and chemical-potential-dependent behavior in QCD.
@@
In each of these holographic models, we computed the butterfly velocity using three distinct methods: entanglement wedge reconstruction, OTOCs, and pole skipping, and found that all three give the same result. This extends previous results by demonstrating that such an equivalence holds in a variety of holographic QCD models as well. In the first three models, we obtained results analytically, while in the fourth model, we obtained results numerically. We further analyzed the behavior of $v_{B}$ with respect to the chemical potential and temperature and observed a universal trend across models. In particular, $v_{B}$ is found to be decreasing/increasing with the chemical potential/temperature. The same result is also true in the simplest holographic charged plasma, i.e., dual to the Reissner-Nordström AdS background (see the Appendix A). Such universal behavior offers potential signatures of quantum chaos in QCD that could be accessible to experimental observation, especially in light of the recently proposed experiments to measure OTOCs in quantum systems. Such universal chaotic features might also provide interesting insights into QCD. Note that $v_{B}$ was also studied recently for non-Abelian gauge theories at high temperatures ($2-10~T_{c}$) using lattice techniques, and it was found that the butterfly velocity does not vary significantly compared to its value at the deconfinement transition temperature [115]. In particular, a variation of about $10\%$ at higher temperatures relative to the deconfinement temperature was observed. This can be compared to the results of Section 4, where the deconfinement temperature can be consistently defined and a variation of $25\%$ is found. Similarly, a variation of $8\%$ is found as we go from $T=150$ to $T=500~\text{MeV}$ in the model of Section 5.
@@
Another important remark that we would like to make is regarding the behavior of $v_{B}$ with respect to the charge. In this work, we primarily focused on studying the variation of $v_{B}$ with respect to the chemical potential, but one can also study the variation with respect to the charge. On analyzing, one finds that $v_{B}$ monotonically decreases even with the charge for all the models presented in our work. It is interesting to note that such a monotonically decreasing trend with charge was also observed in (hyperscaling-violating) Lifshitz theories [59]. Thus, it would be interesting to investigate if this trend is a universal feature of all holographic theories or all quantum systems, in general.
@@
In this work, we restricted our attention to bulk black hole geometries that are isotropic and planar. One can also extend this analysis to black hole geometries that are anisotropic. The work of [116] provides an example to compute butterfly velocity in anisotropic backgrounds. Exploring the effect of magnetic fields on the butterfly velocity in holographic QCD models is another interesting avenue. However, since in most holographic models magnetic field induces an anisotropy, computing the butterfly velocity may not be an easy task. One exception to this is the model discussed in [117]. This model has a $SO(3)$ symmetry in spite of the magnetic field. For this model, one can check, using the methods discussed in this work, that the butterfly velocity monotonically decreases as a function of the magnetic field. It would be interesting to investigate if this decreasing trend is also a feature of other more sophisticated holographic models of magnetized QCD [105, 103].
@@
Before concluding, we would like to make a brief remark on the interplay between the butterfly velocity and phase transitions. In [118], it was conjectured that $v_{B}$ can serve as a probe of phase transitions. That work presented examples within AdS/CMT models supporting this idea. The bottom-up holographic model considered in Section 4 exhibits a Hawking/Page phase transition. Notably, we have plotted $v_{B}$ only for the thermodynamically stable large black hole branch. However, as mentioned earlier, there also exists a thermodynamically unstable small black hole branch, which appears only below a critical chemical potential and vanishes above it. Extending the plot of $v_{B}$ to include this unstable small branch reveals a nonmonotonic behavior for values of the chemical potential below the critical point. In contrast, for values above the critical chemical potential, $v_{B}$ exhibits monotonic behavior once again. This suggests that the onset of nonmonotonicity coincides with the critical chemical potential, providing further evidence in support of the conjecture put forward in [118]. Note that, by taking different profiles of the scale factor $A(z)$, one can also construct gravity solutions exhibiting van der Waals type first-order small/large black hole phase transitions with a second-order critical point [53]. It would be interesting to see if information about such transitions and critical points can also be reflected in the butterfly spectrum. There are already interesting suggestions that such small/large phase transitions can be probed by the Lyapunov exponents [119, 120]. Work in this direction is under progress.
@@ Acknowledgements
Data Availability
The data are available from the authors upon request.
@@
A.1 Background
In this section, we compute $v_{B}$ of the charged plasma corresponding to the Reissner-Nordström AdS (RN-AdS) black hole background. The metric of an RN-AdS black hole in five dimensions is
@@ A.1 Background
A.2 Calculation of $v_{B}$ using entanglement wedge method
Comparison of (A.1) with (2.1) yields the following:
@@ A.2 Calculation of $v_{B}$ using entanglement wedge method
Substituting the horizon values of the metric functions (and their derivatives) into (2.13), we get the following expression for the butterfly velocity:
@@
Figure 10 illustrates the behavior of $v_{B}$ with the RN-AdS background. As shown, the qualitative behavior of $v_{B}$ in this setup closely mirrors the trends observed in the other holographic models discussed in previous sections.
A.3 Calculation of $v_{B}$ using the OTOC method
@@
A.4 Calculation of $v_{B}$ using pole-skipping method
@@
The expression for the butterfly velocity $v_{B}^{2}=\omega_{*}^{2}/q_{*}^{2}$,
$v_{B}^{2}=\frac{\omega^{2}}{q^{2}}=\frac{2}{3}\left(1-\frac{Q^{2}}{2}\right)\,$, (A.11)